SPECTROSCOPY AND DYNAMICS OF THE H_{2}-CN VAN DER WAALS COMPLEX

\underline{YALING CHEN} and MICHAEL C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

The reaction H_{2} + CN → H + HCN is of importance in the combustion of hydrocarbons in air. It is also

a prototypical system for studies of polyatomic reaction dynamics. The reaction has a substantial entrance channel barrier (≈1000cm^{−1}), which raises the possibility that pre-reactive H_{2}-CN van der Waals complex may be stable at low temperatures.

We have recently detected H_{2}-CN in a free-jet expansion, using laser excitation of the B-X and A-X transitions. The complex feature associated with the monomer B-X 0-0 transition consisted of a single broad peak. This corresponds to direct excitation to the H_{2} + CN(B) dissociation continuum. From the onset of the continuum, the ground state well-depth is estimated to be 40cm^{−1}.

Bands of H_{2}-CN associated with the monomer A^{2}Π-X^{2}Σ^{+} 3-0 transition were examined. Complex features

associated with A^{2}Π_{1/2} were found to be homogeneously broadened by the spin-orbit predissociation process

H_{2}-CN(A^{2}Π_{1/2}, v=3) → H_{2} + CN(A^{2}Π_{3/2}, v=3) [Γ=2.5x10^{11}s^{−1}]

Excitation of the complex to the A^{2}Π_{3/2} spin-orbit component yielded rotationally resolved bands. A preliminary analysis of the rotational structure yields an H_{2} to CN separation of 3.7Å. The rotational lines are homogeneously broadened by the internal conversion predissociation process

H_{2}-CN(A^{2}Π_{3/2}, v=3) → H_{2} + CN(X, v=7) [Γ=1.3x10^{10}s^{−1}]

Double resonance techniques have been used to examine the CN fragments produced by both predissociation channels. Data for H_{2}-CN and D_{2}-CN (experiments in progress) will be presented and discussed.